# Forecasting of the time-dependent fluxes of antiprotons in the AMS-02 era

### v1

@@ 1 Introduction
Galactic cosmic rays (GCRs) are charged, energetic particles originating from cosmic accelerators, such as supernova remnants, and propagate throughout the Milky Way (Moskalenko & Strong, 1998). Upon entering the heliosphere, they are modulated by the outward-moving magnetized solar wind plasma (Potgieter, 2013). It is important to study the solar modulation as it is crucial for understanding the nature of GCRs, including their origins (Blasi, 2013) and propagation mechanisms (Strong et al., 2007) within the galaxy. Additionally, solar modulation plays a critical role in the search for dark matter, as it affects low-energy antiproton and antideuteron fluxes (Lavalle & Salati, 2012; Cui et al., 2017; Zhu et al., 2022).
The propagation of GCRs within the heliosphere is described by Parker’s transport equation (TPE) (Parker, 1965). This equation can be solved using various methods, including numerical and analytical approaches. Usually, the force field approximation (FFA) (Gleeson & Axford, 1967, 1968) is used to solve the equation as it is simple and enough to explain the observations. The recent experimental results have achieved significant breakthroughs that are instrumental in understanding the solar modulation effect as well as the GCRs physics, particularly those from Voyager, AMS-02, PAMELA and DAMPE (Stone et al., 2013; Adriani et al., 2011; Aguilar et al., 2017; Ambrosi et al., 2017). To date, only Voyager-1 and Voyager-2 have crossed the heliospheric boundary (Stone et al., 2013, 2019), providing direct measurements of the LIS in the energy range of a few to hundreds of MeV/nucleon. The AMS-02 collaboration has further contributed by publishing high-precision cosmic ray spectra (Aguilar et al., 2021a) and the evolution of some cosmic ray fluxes over time (Aguilar et al., 2018a, b), providing new opportunities to study the solar modulation for cosmic rays (CRs). The FFA cannot explain the fluxes itself as well as the changing over a long period of time with homogeneous parameters (Shen et al., 2021; Long & Wu, 2024). Several methods have been proposed to modify the FFA to account for the cosmic rays fluxes (Corti et al., 2016; Zhu et al., 2021; Shen et al., 2021; Cholis & McKinnon, 2022; Long & Wu, 2024).
@@ 2.4 LIS of Protons, Electrons, Positrons, and Antiprotons
The force-field approximation is applied to derive the LIS for protons, electrons, positrons, and antiprotons. We assume that Voyager data contain only protons or electrons, with no other components, when determining the LIS. Additionally, we assume that protons (antiprotons) and positrons (electrons) share the same solar modulation potential when using AMS-02 data from the same time range Aguilar et al. (2021a). The fitting results are illustrated in Figure 1.
Figure 1: LIS of positron (top left), proton (top right), electron (bottom left) and antiproton (bottom right). The blue points are AMS-02 data from (Aguilar et al., 2021a), and the red points are Voyager data from Cummings et al. (2016). The lines stand for the LIS while the dash lines donate TOA.
@@ 3.2 Prediction of the Time-Dependent Fluxes of Antiprotons
This study is based on the hypothesis that GCRs with the same charge sign exhibit similar solar modulation parameters due to similar diffusion and energy loss processes in the heliosphere. The time-dependent fluxes of protons, helium, and positrons validate this hypothesis for positively charged GCRs. The upcoming AMS-02 time-dependent antiproton flux measurements will provide a critical test for negatively charged GCRs. If confirmed, this finding will profoundly enhance our understanding of the origin and propagation of GCRs in the galaxy. This finding will also be useful for forecasting the time-dependent fluxes of other GCR species, such as Li, Be, B, C, O, and so on.
Thanks for Fan Yi-Zhong, Yuan Qiang and Shu Xin-Wen for very helpful discussions. This work is supported by the National Natural Science Foundation of China (No. 12203103). D.K.K. is support by the National Key Research and Development Program of China (No. 2022YFF0503304). Z.C.R is also support by the Doctoral research start-up funding of Anhui Normal University. We acknowledge the use of data from the AMS Publications  [^2] https://ams02.space/publications/ and the Cosmic-Ray Database (Di Felice et al., 2017)  [^3] https://tools.ssdc.asi.it/CosmicRays/

### v2

@@ 1 Introduction
Galactic cosmic rays (GCRs) are charged, energetic particles originating from cosmic accelerators, such as supernova remnants, and propagate throughout the Milky Way (Moskalenko & Strong, 1998). Upon entering the heliosphere, they are modulated by the outward-moving magnetized solar wind plasma (Potgieter, 2013). It is important to study the solar modulation as it is crucial for understanding the nature of GCRs, including their origins (Blasi, 2013) and propagation mechanisms (Strong et al., 2007) within the galaxy. Additionally, solar modulation plays a critical role in the search for dark matter, as it affects low-energy antiproton and antideuteron fluxes (Lavalle & Salati, 2012; Yuan & Bi, 2015; Cui et al., 2017; Zhu et al., 2022).
The propagation of GCRs within the heliosphere is described by Parker’s transport equation (TPE) (Parker, 1965). This equation can be solved using various methods, including numerical and analytical approaches. Usually, the force field approximation (FFA) (Gleeson & Axford, 1967, 1968) is used to solve the equation as it is simple and enough to explain the observations. The recent experimental results have achieved significant breakthroughs that are instrumental in understanding the solar modulation effect as well as the GCRs physics, particularly those from Voyager, AMS-02, PAMELA and DAMPE (Stone et al., 2013; Adriani et al., 2011; Aguilar et al., 2017; Ambrosi et al., 2017). To date, only Voyager-1 and Voyager-2 have crossed the heliospheric boundary (Stone et al., 2013, 2019), providing direct measurements of the LIS in the energy range of a few to hundreds of MeV/nucleon. The AMS-02 collaboration has further contributed by publishing high-precision cosmic ray spectra (Aguilar et al., 2021a) and the evolution of some cosmic ray fluxes over time (Aguilar et al., 2018a, b), providing new opportunities to study the solar modulation for cosmic rays (CRs). The FFA cannot explain the fluxes itself as well as the changing over a long period of time with homogeneous parameters (Shen et al., 2021; Long & Wu, 2024). Several methods have been proposed to modify the FFA to account for the cosmic rays fluxes (Corti et al., 2016; Yuan et al., 2017; Zhu et al., 2021; Shen et al., 2021; Cholis & McKinnon, 2022; Long & Wu, 2024).
@@ 2.4 LIS of Protons, Electrons, Positrons, and Antiprotons
Based on Voyager data for electrons and protons, we can successfully derive their LISs and corresponding solar modulation potentials using the FFA by assuming that Voyager data contain only protons or electrons, with no other components. Unfortunately, there is no Voyager data for positrons and antiprotons. So, in order to obtain the LISs for positrons and antiprotons, we assume that positrons share the same solar modulation potential as protons and that antiprotons share the same solar modulation potential as electrons, as they have the same charge sign. Since we have obtained the proton LIS in Zhu (2024); Zhu & Wang (2025), we use the same proton LIS here. We assume that positrons share the same solar modulation potential as proton in Zhu (2024) , which is 0.477 GV. Then, we fit the positron LIS while keeping the solar modulation potential unchanged. For the antiproton LIS, we fit it together with the electron LIS and make them share the same solar modulation potential during fitting. Once the LISs are established, they are held constant throughout the analysis. The fitting results are illustrated in Figure 1 and we compare our best-fitting LIS with previous works with somehow different methods and assumptions. Our LISs are very closed to the results from Galprop-HelMod join effort (Boschini et al., 2017, 2018, 2020) for proton, electron and antiproton. The positron LIS used in this work is also close to Vittino et al. (2019) at energy above 2 GeV.
Figure 1: LIS of positron (top left), proton (top right), electron (bottom left) and antiproton (bottom right). The blue points are AMS-02 data from (Aguilar et al., 2021a), and the red points are Voyager data from Cummings et al. (2016). The lines stand for the LIS while the dash lines donate TOA. The black lines correspond to the LISs used in this work while the magenta curves represent reference LIS from literature sources for comparison (Boschini et al., 2017, 2018; Vittino et al., 2019; Boschini et al., 2020).
@@ 3.2 Prediction of the Time-Dependent Fluxes of Antiprotons
During the review process of this work, we noticed that the AMS Collaboration’s published antiproton flux measurements (Aguilar et al., 2025), which coincided with our analysis timeline. Subsequently, we conducted systematic comparisons between our modified FFA model predictions and these benchmark experimental results in Figures 10 and 11. Remarkably, both modified FFA models demonstrate complete consistency with the latest AMS-02 data within the 1 $\sigma$ confidence interval. This empirical validation conclusively demonstrates the validity of our methodological framework in resolving solar modulation of different GCRs.
@@
Figure 10: Zhu’s model predictions for antiproton fluxes (colored lines) are compared with AMS-02 measurements (data points with error bars) (Aguilar et al., 2025).
Figure 11: Long’s model predictions for antiproton fluxes (colored lines) are compared with AMS-02 measurements (data points with error bars) (Aguilar et al., 2025).
@@
This study is based on the hypothesis that GCRs with the same charge sign exhibit similar solar modulation parameters due to similar diffusion and energy loss processes in the heliosphere. The time-dependent fluxes of protons, helium, and positrons validate this hypothesis for positively charged GCRs. The latest AMS-02 time-dependent antiproton flux measurements provide a critical test for negatively charged GCRs and this finding will profoundly enhance our understanding of the origin and propagation of GCRs in the galaxy. This finding will also be useful for forecasting the time-dependent fluxes of other GCR species, such as Ne, Mg, Si and so on.
Thanks for Fan Yi-Zhong, Yuan Qiang and Shu Xin-Wen for very helpful discussions. This work is supported by the National Natural Science Foundation of China (No. 12203103). D.K.K. is supported by the National Key Research and Development Program of China (No. 2022YFF0503304). Z.C.R is also supported by the Doctoral research start-up funding of Anhui Normal University. We acknowledge the use of data from the AMS Publications  [^2] https://ams02.space/publications/ and the Cosmic-Ray Database (Di Felice et al., 2017)  [^3] https://tools.ssdc.asi.it/CosmicRays/